You have 1 m3 of wood and 1 m3 of aluminum. Which is heavier? Prove your answer with calculation.


1
Expert's answer
2022-02-01T21:46:03-0500

Density, "D= m\/v"

Mass of wood = 1× 1500= 1500 kg

Mass of aluminium= 1× 2710= 2710kg

Aluminium is heavier